for a normal distribution with mean 50 and std dev of 5.2 find P90
Answer #1

Given = u = 50 o = 5.2 Here. Za = 20.9 = 1.28 Now, X90 = u +Za* o = 50 +1.28*5.2 = 56.66

Pls Upvote...!!!